# BadBlocks: Low-Cost and Stealthy Backdoor Attacks Tailored for Text-to-Image Diffusion Models

### v1

@@ Abstract
In recent years, Diffusion models have achieved remarkable progress in the field of image generation. However, recent studies have shown that diffusion models are susceptible to backdoor attacks, in which attackers can manipulate the output by injecting covert triggers such as specific visual patterns or textual phrases into the training dataset. Fortunately, with the continuous advancement of defense techniques, defenders have become increasingly capable of identifying and mitigating most backdoor attacks using visual inspection and neural network-based detection methods. However, in this paper, we identify a novel type of backdoor threat that is more lightweight and covert than existing approaches, which we name BadBlocks, requires only about 30% of the computational resources and 20% GPU time typically needed by previous backdoor attacks, yet it successfully injects backdoors and evades the most advanced defense frameworks. BadBlocks enables attackers to selectively contaminate specific blocks within the UNet architecture of diffusion models while maintaining normal functionality in the remaining components. Experimental results demonstrate that BadBlocks achieves a high attack success rate (ASR) and low perceptual quality loss (as measured by FID Score), even under extremely constrained computational resources and GPU time. Moreover, BadBlocks is able to bypass existing defense frameworks, especially the attention-based backdoor detection method, highlighting it as a novel and noteworthy threat. Ablation studies further demonstrate that effective backdoor injection does not require fine-tuning the entire network and highlight the pivotal role of certain neural network layers in backdoor mapping. Overall, BadBlocks significantly reduces the barrier to conducting backdoor attacks in all aspects. It enables attackers to inject backdoors into large-scale diffusion models even using consumer-grade GPUs. Our codes are available at: https://anonymous.4open.science/r/BadBlocks-5FD0.
@@ 1 Introduction
Today, generative models have become some of the most widely adopted models in the field of artificial intelligence, with diffusion models emerging as one of the most effective approaches. Diffusion models are extensively used to generate high-quality images and videos, and they support a wide range of conditional inputs, such as text prompts, reference images, and ControlNet maps to guide the generation process[1, 2].
@@
However, recent studies have shown that diffusion models are vulnerable to backdoor attacks in which attackers can manipulate the output of models by embedding covert triggers in the input[3, 4]. Backdoor attacks pose significant challenges to the safe development of diffusion models. Once the backdoor is activated by attackers, it can lead to unpredictable and potentially harmful outcomes, such as the generation of malicious content or the manipulation of decisions in downstream classifiers[5, 6, 7]. In addition, due to the stealthy nature of backdoors and the diversity of potential triggers, backdoor attacks are widely regarded as one of the most serious threats to diffusion models. In threat scenarios, attackers upload backdoor-injected models to a public platform (e.g., Hugging Face or Github) while falsely claiming it to be benign. When unsuspecting users download and deploy these models, their output typically appears normal and harmless. However, once the attacker activates the backdoor, the target model begins to generate predefined malicious content[8].
@@
The powerful generative capabilities of diffusion models, coupled with their vulnerability to backdoor attacks, have raised widespread concerns. With the increasing release of pre-trained models, users frequently download the necessary models from open platforms according to their specific requirements[9, 10]. However, due to the opacity of black-box models, users often focus solely on whether the pre-trained models can accomplish the intended tasks, while overlooking potential backdoor threats[11, 12, 13]. For example, in classification tasks, attackers can easily manipulate the output by activating backdoors[14]. In face recognition tasks, when facial images contain trigger patterns, the model may misclassify them as any identity predetermined by the attacker[15]. In generative tasks, backdoor attacks often cause the model to produce malicious content, including violent, gory, or otherwise inappropriate imagery, which may further compromise downstream applications[16]. In diffusion models, the objective of backdoor attacks is often to generate a specific image or a particular category of images.
@@
Fortunately, successfully executing covert backdoor attacks remains a highly challenging task. On the one hand, various defense frameworks have been developed to counter backdoor attacks in diffusion models. These frameworks typically achieve backdoor detection and trigger inversion by designing specialized neural network architectures and tailored loss functions[17, 18]. On the other hand, attackers often incur significant costs for computing resources and must invest considerable time in fine-tuning the model to inject covert and effective backdoors, which limits the scalability of backdoor attacks.
@@
  • For the first time, we identify that different layers of neural networks in diffusion models exhibit varying degrees of sensitivity to backdoor attacks. We model the differences among these parameters and highlight the critical role of different components in facilitating backdoor attacks. In addition, we conduct an in-depth analysis of the causes of the "assimilation phenomenon" in Text-to-Image backdoor and reveal the limitations of backdoor attack detection via defense mechanisms based on attention features.
@@
2 Related Works
In this section, we introduce the fundamental principles of diffusion models and review existing backdoor attack techniques applied to them. We also outline previously advanced defense strategies. Finally, we discuss the limitations of both attack and defense techniques.
2.1 Fundamental Principles of Diffusion Models
@@
Diffusion models are widely regarded as one of the most powerful generative models, capable of learning data distributions from the original dataset $x$. Denoising Diffusion Probabilistic Model (DDPM) was the first framework to introduce diffusion models and apply them to image generation tasks[19]. DDPM decomposes the training process of diffusion models into two stages: a forward diffusion process and a reverse denoising process. The forward process is typically represented as $x_{0}\rightarrow x_{t}$. In discrete time steps, $x_{t}$ can be obtained through a noise strength function $\alpha_{t}$ and a noising equation $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\overline{\alpha}}\epsilon$. The reverse process is modeled as a Markov chain of repeated denoising steps, where each step can be represented as: $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, which $\epsilon_{\theta}$ represents the output of the model. Denoising Diffusion Implicit Models (DDIM) significantly accelerate the sampling process by removing the Markovian dependency in the reverse process, thereby reducing the number of sampling steps from 1000 to approximately 100[20]. Latent Diffusion Models (LDMs) compress input data from the pixel space into a latent space using a Variational Autoencoder (VAE), significantly reducing the computational resources required for training and inference[21, 22]. Stochastic Differential Equations (SDE) unifies the forward and reverse processes of the diffusion models into stochastic differential equations and introduces a novel training loss based on the score function $\nabla_{x}\log P_{t}(x)$[23]. By removing stochastic dependency in the reverse process, certain schedulers leverage ordinary differential equations (ODEs) to enable fast sampling, reducing the number of steps required to generate new images to as few as 15 to 30[24, 25].
@@
2.2 Backdoor Attacks in Diffusion Models
Since the introduction of diffusion models, backdoor attacks have been widely recognized as one of their most significant security threats[26]. Attackers can establish covert mappings between neurons by fine-tuning the model, ensuring that specific input vectors yield predefined outputs[27]. TrojDiff was the first method to successfully implement backdoor attacks targeting the noise space in both DDPM and DDIM. Introduced trigger-specific perturbations by injecting crafted noise patterns into the diffusion process, effectively embedding triggers in the noise space[3]. BadDiffusion is conceptually similar to TrojDiff, but it does not rely on trigger embedding functions. Instead, it establishes backdoors by directly injecting patches into the noise space, enabling the generation of specific target images[4]. RickRolling is the first study to extend the attack space to the textual input space, enabling backdoor attacks on Latent Diffusion Models (LDMs), particularly Stable Diffusion[28]. Furthermore, VillanDiffusion introduced a backdoor attack framework that unifies text and noise attack spaces, allowing a wider range of attack targets[29]. Subsequently, additional attack spaces were explored. For example, Gungnir employed the style of the input image as a backdoor trigger in image-to-image tasks[30]. It is evident that attackers are continuously expanding the backdoor attack space, which undoubtedly presents a significant challenge to the security of diffusion models.
2.3 Defense Strategies in Diffusion Models
As an increasing number of backdoor threats are revealed, a growing body of defense techniques against backdoor attacks on diffusion models has been proposed. Elijah is the first defense framework specifically designed to defend against backdoor attacks in diffusion models and has successfully mitigated attacks from BadDiff, TrojDiff, and VillanDiffusion[17]. TERD builds upon Elijah by unifying the attack formulation within the noise attack space as $x_{t}=a(x_{0},t)x_{0}+b(t)\epsilon+c(t)r$. Among them, $\alpha$ and $b$ represent the original image-noise ratio function, and $c$ is the trigger injection function. TERD achieves high precision backdoor detection and trigger inversion by computing the KL divergence between inverted and benign noise distributions. For the text-based attack space, T2IShield is the first defense framework proposed to detect backdoor attacks triggered by prompt words[31]. By analyzing the attention maps during the diffusion process and identifying the "Assimilation Phenomenon", T2IShield designs a defense algorithm based on attention maps, effectively defending against attacks such as RickRolling. As more covert and efficient backdoor attack methods continue to emerge, defense strategies are also evolving, from the early stages of designing specific neural networks and loss functions to more recent efforts that analyze distributional differences within black-box models. It is evident that these defense mechanisms are becoming increasingly fine-grained. However, a vast attack space remains unexplored, including areas such as ControlNet and adapter-based tasks[32].
2.4 Limitations
From the perspective of attack strategies, injecting backdoors into diffusion models faces two major challenges: (1) it often demands substantial computational resources, especially GPU memory, as attack diffusion models can be as costly as full-scale training; and (2) target-specific poisoning commonly results in the "assimilation phenomenon," which increases the risk of detection by defense mechanisms. From the defense perspective, current methods lack adaptability across different attack spaces. For instance, Elijah and TERD are constrained to detecting backdoors originating in the noise space, while T2IShield is only effective against prompt-based attacks.
3 Method
In this section, we introduce the BadBlocks threat model, outlining its attack scenarios and background. We analyze the capabilities of both attackers and defenders, and briefly describe the attack method with a theoretical explanation of its feasibility.
3.1 Threat Model
The threat model for backdoor attacks on diffusion models is typically defined as a black-box setting from the user’s perspective[33, 34]. Depending on the attack method, adversaries can poison the training data or alter the training strategy, ultimately providing backdoored models to the user[35]. The threat model in BadBlocks is largely consistent with previous work. However, because the weight distribution and features of the remaining blocks remain unchanged, BadBlocks is more likely to evade detection by defense frameworks, especially which methods based on the global features of the UNet structure. You can find the analysis of the differences in the distribution of model weights in Appendix.A and B.
3.2 Attacker and Defender Knowledge
Building on existing work[36, 37], we define the knowledge of the attacker in the BadBlocks setting as follows: (1) The attacker has the authority to manipulate the fine-tuning process of the model, including the dataset, the loss function and the training strategy; (2) The attacker can access the model and provide input data that comply with its specifications. In contrast, a model defender not only has full access to all model parameters and can input arbitrary data to obtain detection samples, but can also perform additional operations such as distillation, pruning, and other modifications about model weights [38, 39].
3.3 Mitigation Assimilation
@@
Inspired by previous studies[40, 41], we recognize that different neural network blocks within the UNet architecture of a diffusion model serve distinct roles during the image generation process. Therefore, we attempt to design an algorithm that decouples the neural components associated with backdoor mapping from the UNet and applies fine-grained training to them. Specifically, in the attention-based detection method, given a token sequence of length $L$, each block $b$ generates a cross-attention map $M_{b}^{(t)}$ at any time step $t\in T$:
@@
However, in BadBlocks, the attention maps $M_{b}$ generated by most blocks appear to be entirely normal, as their parameters remain unchanged. Only the injected sampled blocks that are targeted by the attack exhibit abnormal attention patterns, and these anomalies can be easily masked when averaged with the attention maps from benign blocks. This allows the remaining benign cross-attention layers to suppress the assimilation effect induced by the backdoor. Ultimately, through controlled experiments, we successfully isolated the network blocks essential for executing backdoor attacks in diffusion models, specifically, network blocks in the upsampling stage and its critical subcomponents: the the resnets[42], the transformer blocks[43] and the normalization layers[44] among them.
Method Infected Components Affected Parameters Necessary GPU Memory GPU Time(1 epcoh)
@@
Table 1: We evaluated the minimum computational overhead of BadBlocks for 20,000 samples on an NVIDIA A40 GPU. Compared to existing methods, BadBlocks introduces significantly less impact on the target model. Moreover, its low dependency on attacker-side hardware substantially lowers the barrier to executing backdoor attacks.
@@ 3.4 Approaches of BadBlocks
We adopt the definition of the input space $S_{input}$ from Gungnir[30], where the additional conditions space $a\subseteq{A_{cond}}$ is introduced to constrain the stochasticity of the diffusion process. The complete composition of the additional space $A_{cond}$ is:
@@
In BadBlocks, we define a hyperparameter $i\geq 0$ to represent the number of blocks involved in backdoor training, where $i$ is strictly less than the total number of upsample blocks. Importantly, during training, we froze nearly all parameters of the model $M_{\theta}$, performing gradient updates only on the critical weight $\theta^{*}_{i}$, including the ResNet layers, Transformer blocks, and normalization layers. Figure.3 illustrates the specific parameters within the sampled blocks targeted by BadBlocks that require training.
@@
where $k$ is the total number of parameter tensors in the block. In the experimental section, we leverage this differential analysis to identify the parameters that play a pivotal role in establishing backdoor mappings within neural networks, thereby providing critical evidence to support our findings. Figure.2 visualizes the differences in weight changes across components affected by BadBlocks.
4 Experiments
@@
In this section, we present the evaluation results of BadBlocks in the text-to-image task, along with its corresponding attack settings. Figure.4 shows the attack process of BadBlocks. Furthermore, we highlight the advantages of our method in terms of training efficiency and resource consumption while also acknowledging its limitations.
4.1 Attack Configuration

### v4

# BadBlocks: Lightweight and Stealthy Backdoor Threat in Text-to-Image Diffusion Models

@@ Abstract
Diffusion models have recently achieved remarkable success in image generation, yet growing evidence shows their vulnerability to backdoor attacks, where adversaries implant covert triggers to manipulate outputs. While existing defenses can detect many such attacks via visual inspection and neural network–based analysis, we identify a more lightweight and stealthy threat, termed BadBlocks. BadBlocks selectively contaminates specific blocks within the UNet architecture while preserving the normal behavior of the remaining components. Compared with prior methods, it requires only about 30% of the computation and 20% of the GPU time, yet achieves high attack success rates with minimal perceptual degradation. Extensive experiments demonstrate that BadBlocks can effectively evade state-of-the-art defenses, particularly attention-based detection frameworks. Ablation studies further reveal that effective backdoor injection does not require fine-tuning the entire network and highlight the critical role of certain layers in backdoor mapping. Overall, BadBlocks substantially lowers the barrier for backdooring large-scale diffusion models, even on consumer-grade GPUs.
@@ I Introduction
Today, generative models have become some of the most widely adopted models in the field of artificial intelligence, with diffusion models emerging as one of the most effective approaches. Diffusion models are extensively used to generate high-quality images and videos, and they support a wide range of conditional inputs, such as text prompts, reference images, and ControlNet maps to guide the generation process[diffusion_survey1, diffusion_survey2].
@@
However, recent studies have shown that diffusion models are vulnerable to backdoor attacks in which attackers can manipulate the output of models by embedding covert triggers in the input[trojdiff, baddiffusion]. Backdoor attacks pose significant challenges to the safe development of diffusion models. Once the backdoor is activated by attackers, it can lead to unpredictable and potentially harmful outcomes, such as the generation of malicious content or the manipulation of decisions in downstream classifiers[survey, survey2, survey3]. In addition, due to the stealthy nature of backdoors and the diversity of potential triggers, backdoor attacks are widely regarded as one of the most serious threats to diffusion models. In threat scenarios, attackers upload backdoor-injected models to a public platform (e.g., Hugging Face or Github) while falsely claiming it to be benign. When unsuspecting users download and deploy these models, their output typically appears normal and harmless. However, once the attacker activates the backdoor, the target model begins to generate predefined malicious content[BadT2I].
@@
The powerful generative capabilities of diffusion models, coupled with their vulnerability to backdoor attacks, have raised widespread concerns. With the increasing release of pre-trained models, users frequently download the necessary models from open platforms according to their specific requirements[pretrained:1, pretrained:2]. However, due to the opacity of black-box models, users often focus solely on whether the pre-trained models can accomplish the intended tasks, while overlooking potential backdoor threats[backdoor:1, backdoor:2, backdoor:3]. For example, in classification tasks, attackers can easily manipulate the output by activating backdoors[classback]. In face recognition tasks, when facial images contain trigger patterns, the model may misclassify them as any identity predetermined by the attacker[faceback]. In generative tasks, backdoor attacks often cause the model to produce malicious content, including violent, gory, or otherwise inappropriate imagery, which may further compromise downstream applications[llmback]. In diffusion models, the objective of backdoor attacks is often to generate a specific image or a particular category of images.
@@
Fortunately, successfully executing covert backdoor attacks remains a highly challenging task. On the one hand, various defense frameworks have been developed to counter backdoor attacks in diffusion models. These frameworks typically achieve backdoor detection and trigger inversion by designing specialized neural network architectures and tailored loss functions[eliagh, terd]. On the other hand, attackers often incur significant costs for computing resources and must invest considerable time in fine-tuning the model to inject covert and effective backdoors, which limits the scalability of backdoor attacks.
@@
  • For the first time, we identify that different layers of neural networks in diffusion models exhibit varying degrees of sensitivity to backdoor attacks. We model the differences among these parameters and highlight the critical role of different components in facilitating backdoor attacks. In addition, we conduct an in-depth analysis of the causes of the ”assimilation phenomenon” in Text-to-Image backdoor and reveal the limitations of backdoor attack detection via defense mechanisms based on attention features.
@@
II Related Works
This section reviews the fundamentals of diffusion models, existing backdoor attacks, and representative defense strategies, and summarizes their limitations.
II-A Fundamental Principles of Diffusion Models
@@
Diffusion models are powerful generative models that learn data distributions from training samples $x$. Denoising Diffusion Probabilistic Models (DDPM) first formalized diffusion-based image generation by modeling a forward noising process $x_{0}\rightarrow x_{t}$ and a reverse denoising process[ddpm]. The forward process follows $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\overline{\alpha}}\epsilon$, while the reverse step is given by $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\epsilon,\ \epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\epsilon_{\theta}$ is the model prediction. Denoising Diffusion Implicit Models (DDIM) remove the Markovian dependency to accelerate sampling[ddim]. Latent Diffusion Models (LDMs) compress data into a latent space using a VAE to reduce computational cost[stablediffusion, vae]. Stochastic Differential Equations (SDE) unify diffusion processes under continuous formulations and introduce score-based training[sde]. Further, ODE-based schedulers enable fast sampling within 15–30 steps[dpmv1, dpmv3].
II-B Backdoor Attacks in Diffusion Models
Backdoor attacks have emerged as a critical threat to diffusion models[riskindiff]. By fine-tuning models, attackers establish covert neuron-level mappings to force predefined outputs[backdoor:4]. TrojDiff first injected trigger-specific perturbations into the noise space of DDPM and DDIM[trojdiff], while BadDiffusion directly embedded patches into the noise space without trigger functions[baddiffusion]. RickRolling extended attacks to the textual prompt space in LDMs, especially Stable Diffusion[rickroll]. VillanDiffusion unified noise and text attack spaces for broader targets[villan]. Beyond these, Gungnir exploited image style as triggers in image-to-image tasks[gungnir]. Overall, the backdoor attack space is rapidly expanding, posing increasing security risks.
II-C Defense Strategies in Diffusion Models
To counter these threats, various defenses have been proposed. Elijah is the first framework tailored for diffusion models, mitigating attacks such as BadDiff, TrojDiff, and VillanDiffusion[eliagh]. TERD extends Elijah by modeling noise-space attacks as $x_{t}=a(x_{0},t)x_{0}+b(t)\epsilon+c(t)r$, and detects backdoors via KL divergence between benign and inverted noise distributions. For text-based attacks, T2IShield analyzes attention maps and the ”Assimilation Phenomenon” to defend against prompt-triggered backdoors like RickRolling[T2IShield]. Recent defenses increasingly rely on distributional analysis within black-box models rather than handcrafted architectures. However, emerging scenarios such as ControlNet and adapter-based tasks remain underexplored[IP-Adapter].
II-D Limitations
Existing attacks face two main challenges: (1) high computational cost comparable to full model training, and (2) target-specific poisoning often induces the ”assimilation phenomenon,” increasing detectability. From the defense perspective, current methods lack generalization across attack spaces: Elijah and TERD focus on noise-based attacks, whereas T2IShield is limited to prompt-based scenarios.
III Method
In this section, we introduce the BadBlocks threat model, outlining its attack scenarios and background. We analyze the capabilities of both attackers and defenders, then briefly describe the attack method with a theoretical explanation of its feasibility.
III-A Threat Model
The threat model for backdoor attacks on diffusion models is typically defined as a black-box setting from the user’s perspective[bagm, picc]. Depending on the attack method, adversaries can poison the training data or alter the training strategy, ultimately providing backdoored models to the user[dpf]. The threat model in BadBlocks is largely consistent with previous work. However, because the weight distribution and features of the remaining blocks remain unchanged, BadBlocks is more likely to evade detection by defense frameworks, especially which methods based on the global features of the UNet structure. You can find the analysis of the differences in the distribution of model weights in Appendix A and B. Building on existing work[BadNL, DBA], we define the knowledge of the attacker in the BadBlocks setting as follows: (1) The attacker has the authority to manipulate the fine-tuning process of the model, including the dataset, the loss function and the training strategy; (2) The attacker can access the model and provide input data that comply with its specifications. In contrast, a model defender not only has full access to all model parameters and can input arbitrary data to obtain detection samples, but can also perform additional operations such as distillation, pruning, and other modifications about model weights [distillerase, diff-clean].
@@ III-A Threat Model
III-B Mitigation Assimilation
@@
Inspired by previous studies[hm, le], we recognize that different neural network blocks within the UNet architecture of a diffusion model serve distinct roles during the image generation process. Therefore, we attempt to design an algorithm that decouples the neural components associated with backdoor mapping from the UNet and applies fine-grained training to them. Specifically, in the attention-based detection method, given a token sequence of length $L$, each block $b$ generates a cross-attention map $M_{b}^{(t)}$ at any time step $t\in T$:
@@
However, in BadBlocks, the attention maps $M_{b}$ generated by most blocks appear to be entirely normal, as their parameters remain unchanged. Only the injected sampled blocks that are targeted by the attack exhibit abnormal attention patterns, and these anomalies can be easily masked when averaged with the attention maps from benign blocks. This allows the remaining benign cross-attention layers to suppress the assimilation effect induced by the backdoor. Ultimately, through controlled experiments, we successfully isolated the network blocks essential for executing backdoor attacks in diffusion models, specifically, network blocks in the upsampling stage and its critical subcomponents: the the resnets[resnet], the transformer blocks[trans] and the normalization layers[groupnorm] among them.
TABLE I: We evaluated the minimum computational overhead of BadBlocks for 20,000 samples on an NVIDIA A40 GPU. Compared to existing methods, BadBlocks introduces significantly less impact on the target model. Moreover, its low dependency on attacker-side hardware substantially lowers the barrier to executing backdoor attacks.
Method Infected Components Affected Parameters Necessary GPU Memory GPU Time (1 epoch)
@@ III-C Approaches of BadBlocks
We adopt the definition of the input space $S_{input}$ from Gungnir[gungnir], where the additional conditions space $a\subseteq{A_{cond}}$ is introduced to constrain the stochasticity of the diffusion process. The complete composition of the additional space $A_{cond}$ is:
@@
In BadBlocks, we define a hyperparameter $i\geq 0$ to represent the number of blocks involved in backdoor training, where $i$ is strictly less than the total number of upsample blocks. Importantly, during training, we froze nearly all parameters of the model $M_{\theta}$, performing gradient updates only on the critical weight $\theta^{*}_{i}$, including the ResNet layers, Transformer blocks, and normalization layers. Figure reffig:fig4 illustrates the specific parameters within the sampled blocks targeted by BadBlocks that require training.
@@
where $k$ is the total number of parameter tensors in the block. In the experimental section, we leverage this differential analysis to identify the parameters that play a pivotal role in establishing backdoor mappings within neural networks, thereby providing critical evidence to support our findings. Figure 10 visualizes the differences in weight changes across components affected by BadBlocks.
IV Experiments
@@
In this section, we present the evaluation results of BadBlocks in the text-to-image task, along with its corresponding attack settings. Figure 4 shows the attack process of BadBlocks. Furthermore, we highlight the advantages of our method in terms of training efficiency and resource consumption while also acknowledging its limitations.
IV-A Attack Configuration